OpenStudy (anonymous):

what is the domain of the function? y=√5x-10 (btw the root covers 5x-10)

jagr2713 (jagr2713):

i got the same answer as the question

OpenStudy (isaiah.feynman):

Good thing you specified. :)

OpenStudy (isaiah.feynman):

Now square cannot contain negative numbers. Therefore the expression in that square root must be GREATER THAN OR EQUAL TO ZERO.

OpenStudy (isaiah.feynman):

root*

OpenStudy (isaiah.feynman):

\[\sqrt{5x-10}\ge0\]

OpenStudy (isaiah.feynman):

Solve for x.

OpenStudy (anonymous):

No squareroot

OpenStudy (isaiah.feynman):

Yes.

OpenStudy (anonymous):

\[5x-10\ge0\]

OpenStudy (isaiah.feynman):

I'm multitasking here!

OpenStudy (isaiah.feynman):

So solve for x and determine the domain.

OpenStudy (isaiah.feynman):

\[x \ge 2\] or \[[2,\infty)\]

OpenStudy (isaiah.feynman):

This means that numbers less than 2 are not allowed in that function!

OpenStudy (anonymous):

so the answer is 

OpenStudy (anonymous):

????? @Isaiah.Feynman

OpenStudy (anonymous):

is that the answer? @Isaiah.Feynman